# Time-Scale Coupling Between States and Parameters in Recurrent Neural Networks

### v1

@@ Abstract
We study how gating mechanisms in recurrent neural networks (RNNs) implicitly induce adaptive learning-rate behavior, even when training is carried out with a fixed, global learning rate. This effect arises from the coupling between state-space time scales–parametrized by the gates–and parameter-space dynamics during gradient descent. By deriving exact Jacobians for leaky-integrator and gated RNNs, we obtain a first-order expansion that makes explicit how constant, scalar, and multi-dimensional gates reshape gradient propagation, modulate effective step sizes, and introduce anisotropy in parameter updates. These findings reveal that gates not only control memory retention in the hidden states, but also act as data-driven preconditioners that adapt optimization trajectories in parameter space. We further draw formal analogies with learning-rate schedules, momentum, and adaptive methods such as Adam, showing that these optimization behaviors emerge naturally from gating. Numerical experiments confirm the validity of our perturbative analysis, supporting the view that gate-induced corrections remain small while exerting systematic effects on training dynamics. Overall, this work provides a unified dynamical-systems perspective on how gating couples state evolution with parameter updates, explaining why gated architectures achieve robust trainability and stability in practice.
@@ 1 Introduction
Recurrent neural networks (RNNs) remain a fundamental tool for modeling sequential data, yet their training dynamics continue to present theoretical and practical challenges. While much progress has been made in designing architectures that mitigate gradient pathologies, improve stability, or enhance memory retention, a unified understanding of how state dynamics and parameter updates interact—particularly under the influence of gating—remains incomplete. In many modern architectures, gates act not only as information filters but also as modulators of time scales in the underlying dynamics. This dual role raises important questions about how gating shapes both the short-term evolution of hidden states and the long-term trajectory of parameters during optimization.
@@
In this work, we show that gating mechanisms have a further, often overlooked effect: they implicitly induce adaptive learning-rate behavior, even when training is performed with a static, fixed global learning rate. This phenomenon arises because the same time-scale modulations that gates impose on state dynamics directly couple into the parameter-update dynamics via backpropagation. In other words, the gates do not merely control information flow in the hidden states; they also reshape the effective optimization process in parameter space, altering step sizes, scaling directions, and introducing anisotropy in the updates.
@@
Our analysis is grounded in a perturbative expansion of the Jacobian products arising in backpropagation, which makes these effects explicit. We derive exact Jacobians for leaky-integrator, scalar-gated, and multi-gated RNNs, and show how their influence on state transitions translates into systematic modifications of parameter-space dynamics. This perspective reveals clear analogies with gradient descent algorithms such as SGD with learning rate schedules, momentum, and adaptive methods like Adam. Importantly, these behaviors emerge endogenously from the gating dynamics, rather than from externally imposed optimization heuristics.
Beyond theoretical analysis, we validate our assumptions and approximations through numerical experiments on RNNs with different gating configurations. These experiments confirm the accuracy of the first-order Jacobian expansion in both small-perturbation and practical operating regimes, supporting the conclusion that gates exert structured but bounded corrections to the dominant state-transition dynamics. Taken together, our findings provide a unified dynamical-systems perspective on how gating couples state evolution with parameter updates, explaining why gated architectures achieve robust trainability and stability in practice.
The main contributions of this paper are:
  • We present a unified theoretical framework that explains how gates implicitly induce adaptive learning-rate effects in parameter space, even under a fixed global learning rate, by coupling state-space time scales with parameter dynamics.
  • We derive explicit Jacobian-based perturbative expansions for constant, scalar, and multi-dimensional gates, linking their influence to gradient propagation, anisotropy in parameter updates, and memory retention.
  • We establish formal analogies between gating-induced effects and optimization algorithms, including learning-rate schedules, momentum, and adaptive methods such as Adam.
  • We validate our theoretical framework with numerical experiments, confirming both the smallness of gate-induced corrections and the robustness of the first-order approximation.
@@
The remainder of the paper is organized as follows. Section 2 surveys the most relevant literature on gradient pathologies, linear and state-space models, constrained recurrent dynamics, and gating mechanisms. Section 3 introduces the interpretation of RNN gating mechanisms as parametrized time-scales and formalizes the specific models considered in our analysis. Section 4 reviews backpropagation through time in RNNs and revisits the vanishing/exploding gradient problem in this context. Section 5 presents our main contribution: we derive the Jacobian matrices for the considered RNN models and show how state-space time-scales influence the behaviour of gradient descent, thereby revealing the interaction between state and parameter dynamics. Section 6 discusses well-known variants of gradient descent and establishes explicit connections between their behaviour and the effects induced by gating mechanisms. Finally, Section 7 summarizes the key findings and outlines future research directions, including the extension of our framework to more advanced gated architectures such as LSTMs and GRUs.
@@ 3.3 General Time Warping and Gating
where $\odot$ denotes elementwise multiplication and
$g_{t}=\begin{bmatrix}g_{t}^{(1)}\\ \vdots\\ g_{t}^{(N_{r})}\end{bmatrix}\in(0,1)^{N_{r}}.$ (17)
Here $\sigma(\cdot)$ in (14) is applied componentwise, with $W^{r,g}\in\mathbb{R}^{N_{r}\times N_{r}}$ and $W^{i,g}\in\mathbb{R}^{N_{r}\times N_{i}}$. Each neuron thus possesses its own time-dependent update rate (or time scale), enabling fine-grained adaptation of the network’s temporal dynamics.
@@ 4 BPTT-based training of RNNs
4.1 Gradient descent in neural networks
The most basic form of the gradient descent (GD) update rule for a parameter vector $\theta$ using a mini-batch of size $b\geq 1$ is
@@ 4.1 Gradient descent in neural networks
where $l$ indexes the training iteration, $\mu>0$ is the learning rate (or step size), and $\mathcal{E}_{t}$ is the loss at sample $t$ in the mini-batch. For instance, in a regression setting one may use the squared error
@@
where $y_{t}$ is the target output, $z_{t}=f(\cdot;\theta)$ is the model prediction (combining the state-update and output mappings), and $\theta$ collects all trainable parameters of the network.
@@ 5.2 Time-scale interaction
In this section, we show that the interaction between state-space time scales and parameter updates gives rise to an effective learning rate $\mu^{*}$, which generally differs from the nominal learning rate $\mu$ in (18). The precise form of $\mu^{*}$ depends on the chosen state-space model. To this end, without loss of generality we consider only time-step $t$ and expand the gradient formula in (19) as follows:
@@ 5.2.1 Constant gate case
$\mu^{*}_{t,k}=\mu\,\alpha^{\,t-k}.$ (44)
This makes explicit that the learning rate decays exponentially with temporal distance, at a rate fully determined by $\alpha$. For $\alpha<1$, long-range dependencies are progressively downweighted, leading to a vanishing gradient phenomenon for long temporal distances.
@@ 5.2.2 RNNs with scalar gate
In what follows and in Section 5.2.3 we use the general first-order expansion described in Appendix A applied to the product of Jacobian matrices with $\epsilon=1$.
@@
where $A_{j-1}=D_{j-1}W^{r}-I$ and $G_{j-1}=d_{j-1}(J^{g}_{j-1})$ is a rank-1 matrix resulting from the outer product of two vectors.
Expanding the product to the first order in the rank-1 corrections $G_{j-1}$, we obtain
@@
The first term describes the main dynamics without the rank-1 updates, while the second term is a sum of rank-1 corrections, each inserted at a different time step $m$ and propagated by the full-rank dynamics. We now factor the gate scalars $g_{j-1}$ explicitly from the first term in (48) like we did in Section 5.2.1.
@@
$\mu^{*}_{t,k}=\mu\prod_{j=k+1}^{t}g_{j-1},$ (51)
where $\mu^{*}_{t,k}$ depends on the entire input-driven trajectory of the gates. The additional rank-1 terms capture further corrections due to the gating mechanism, modulating the propagation of gradients in a way that cannot be simplified into a single scalar factor.
@@ 5.2.3 RNNs with multiple gates
The first term describes the main dynamics through diagonal gates and normalized matrices $\tilde{A}_{j-1}$. The second term accounts for full-rank corrections due to $G_{j-1}$.
@@
$\mu^{*(i)}_{t,k}=\mu\prod_{j=k+1}^{t}g^{(i)}_{j-1}$, (59)
where $\mu^{*(i)}_{t,k}$ is the effective learning rate for neuron $i$. This shows that the gating mechanism induces an anisotropic, path-dependent scaling of the gradient. For the multi-gate case:
  • Each neuron has its own time-varying effective learning rate determined by its gate trajectory.
  • The propagation of gradients is anisotropic, with different attenuation along different state-space directions.
  • The correction terms involving $G_{j-1}$ are full-rank and cannot be simplified into a low-dimensional form, although their effect is modulated by products of diagonal gate matrices.
Therefore, gating with neuron-specific time scales modulates both the magnitude and directionality of gradient propagation, offering a powerful mechanism for controlling learning dynamics.
6 Gradient descent algorithms
@@
6.1 Adam
$\hat{m}_{t,l}$ $=\frac{m_{t,l}}{1-\beta_{1}^{l}},$ (62)
$\hat{v}_{t,l}$ $=\frac{v_{t,l}}{1-\beta_{2}^{l}}.$ (63)
The parameter update is then:
6.2 Analogies between gating mechanisms and adaptive gradient descent methods
@@
With a scalar gate $g_{j-1}$, the effective learning rate becomes
$\mu^{*}_{t,k}=\mu\prod_{j=k+1}^{t}g_{j-1}.$
Here, the gradient is modulated by a global, input-driven attenuation factor that varies over time. This is conceptually similar to a learning rate schedule (e.g., exponential decay), with the key difference that the modulation emerges from the network’s own state dynamics rather than from an externally prescribed schedule.
@@
In the multi-gate case, the modulation is neuron-specific:
$\mu^{*(i)}_{t,k}=\mu\prod_{j=k+1}^{t}g^{(i)}_{j-1}$,
so each neuron $i$ has its own effective learning rate, determined by the trajectory of its gate values. This mirrors adaptive optimizers such as Adam or RMSProp, which assign each parameter a distinct, dynamically adjusted step size.

### v2

@@ Abstract
We study how gating mechanisms in recurrent neural networks (RNNs) implicitly induce adaptive learning-rate behavior, even when training is carried out with a fixed, global learning rate. This effect arises from the coupling between state-space time scales–parametrized by the gates–and parameter-space dynamics during gradient descent. By deriving exact Jacobians for leaky-integrator and gated RNNs, we obtain a first-order expansion that makes explicit how constant, scalar, and multi-dimensional gates reshape gradient propagation, modulate effective step sizes, and introduce anisotropy in parameter updates. These findings reveal that gates not only control information flow, but also act as data-driven preconditioners that adapt optimization trajectories in parameter space. We further draw formal analogies with learning-rate schedules, momentum, and adaptive methods such as Adam, pointing to possible redundancies. Empirical simulations corroborate these claims: in canonical synthetic sequence tasks (adding, copy) we show that gates induce lag-dependent effective learning rates and directional concentration of gradient flow, with multi-gate models matching or exceeding the anisotropic structure produced by Adam. These results highlight that optimizer-driven and gate-driven adaptivity are complementary but not equivalent mechanisms. Overall, this work provides a unified dynamical-systems perspective on how gating couples state evolution with parameter updates, explaining why gated architectures achieve robust trainability and stability in practice.
@@ 1 Introduction
Recurrent neural networks (RNNs) remain a central tool for modeling sequential data, yet their training dynamics continue to raise theoretical and practical challenges. A large body of work has shown how gating mechanisms—as in LSTMs and GRUs—help stabilize learning by regulating information flow in hidden states and mitigating gradient pathologies. From this perspective, gates are primarily understood as filters that control memory retention and short-term state evolution.
@@
In this work, we demonstrate that gates have an additional, previously overlooked role: they also reshape the parameter dynamics during optimization. Specifically, we show that gating mechanisms implicitly induce adaptive learning-rate behavior, even when training is carried out with a fixed global step size. The same time-scale modulations that gates impose on state dynamics couple into parameter updates via backpropagation, altering effective step sizes, introducing lag dependence, and concentrating gradient flow into dominant directions. Thus, gates act not only as information filters but also as data-driven preconditioners of the optimization process.
@@
Our theoretical analysis is based on a perturbative expansion of the Jacobian products arising in backpropagation. By deriving exact Jacobians for leaky-integrator, scalar-gated, and multi-gated RNNs, we make explicit how gates renormalize effective learning rates and introduce anisotropy in parameter updates. This perspective reveals structural analogies with optimization heuristics such as learning-rate schedules, momentum, and adaptive methods like Adam—yet these behaviors emerge endogenously from gating rather than being externally imposed.
We complement this analysis with controlled simulations that probe the predicted phenomena. In Section 7.1, we quantify how gates modulate lag-dependent effective learning rate profiles. In Section 7.2, we introduce an anisotropy index to measure how gates and optimizers shape the directional structure of gradient propagation. These experiments confirm that gate-induced corrections remain bounded but systematic, providing quantitative evidence that gates influence both the magnitude and orientation of temporal credit assignment. Section 7.3 discusses the broader implications of these findings for the trainability and stability of gated architectures.
Contributions.
We propose a unified dynamical-systems framework for analyzing how gating mechanisms affect the coupled dynamics of states and parameters in RNNs. Our contributions are:
  • We show analytically that gates act as parametrized time scales that modulate lag-dependent effective learning rates.
  • We extend the analysis to directional effects, introducing the anisotropy index to quantify how gates and optimizers shape the dominant subspaces of gradient propagation.
  • We validate these predictions through targeted simulations on canonical tasks (adding, copy), comparing leaky, scalar-gated, and multi-gated RNNs as well as adaptive optimization (Adam).
Together, these results highlight that both gating and optimizer dynamics shape not only the magnitude but also the directional structure of temporal credit assignment, with task-dependent trade-offs.
Structure.
@@
The remainder of the paper is organized as follows. Section 2 surveys the most relevant literature on gradient pathologies, linear and state-space models, constrained recurrent dynamics, and gating mechanisms. Section 3 introduces the interpretation of RNN gating mechanisms as parametrized time scales and formalizes the specific models considered in our analysis. Section 4 reviews backpropagation through time in RNNs and revisits the vanishing/exploding gradient problem in this context. Section 5 presents our main theoretical contribution: the derivation of Jacobian matrices for the considered RNN models and the link between state-space time scales and parameter-space dynamics. Section 6 discusses well-known variants of gradient descent and suggests explicit connections between their behaviour and the effects induced by gating mechanisms. Section 7 complements the theory with simulations on well-known sequence tasks. Section 8 summarizes the key insights and outlines future research directions. The paper contains an Appendix with two sections. The first one, Appendix A describes the first-order expansion technique used in this paper and provides numerical examples justifying its robustness. The second one, Appendix B provides a description of the well-know Adam algorithm.
@@ 3.3 General time warping and gating
where $\odot$ denotes elementwise multiplication and $g_{t}\in(0,1)^{N_{r}}$. Here $\sigma(\cdot)$ in (14) is applied componentwise, with $W^{r,g}\in\mathbb{R}^{N_{r}\times N_{r}}$ and $W^{i,g}\in\mathbb{R}^{N_{r}\times N_{i}}$. Each neuron thus possesses its own time-dependent update rate (or time scale), enabling fine-grained adaptation of the network’s temporal dynamics.
@@ 4 BPTT-based training of RNNs
4.1 Stochastic gradient descent in neural networks
The most basic form of the stochastic gradient descent (SGD) update rule for a parameter vector $\theta$ using a randomly sampled mini-batch of size $b\geq 1$ is
@@ 4.1 Stochastic gradient descent in neural networks
where $l$ indexes the training iteration, $\mu>0$ is the learning rate (or step size), and $\mathcal{E}_{t}$ is the loss at sample $t$ in the mini-batch. The stochasticity arises because each mini-batch is drawn at random from the training set, leading to noisy gradient estimates. For instance, in a regression setting one may use the squared error
@@
where $y_{t}$ is the target output, $z_{t}$ is the model prediction, and $\theta$ collects all trainable parameters of the network.
@@ 5.2 Time-scale interaction
In this section, we show that the interaction between state-space time scales and parameter updates gives rise to an effective learning rate $\mu^{*}$, which generally differs from the nominal learning rate $\mu$ in (17). The precise form of $\mu^{*}$ depends on the chosen state-space model. To this end, without loss of generality we consider only one time-step $t$ and expand the gradient formula in (18) as follows:
@@ 5.2.1 Constant gate case
$\mu^{*}_{t,k}\;=\;\mu\,\alpha^{\,t-k}\;+\;\text{(perturbative correction terms)}.$ (43)
Here the perturbative correction terms stem from the residual $(1-\alpha)I$ contribution and from activation derivatives, which slightly distort the pure exponential decay. This makes explicit that the learning rate decays exponentially with temporal distance, at a rate mainly determined by $\alpha$. For $\alpha<1$, long-range dependencies are progressively downweighted, leading to a vanishing gradient phenomenon for long temporal distances.
@@ 5.2.2 RNNs with scalar gate
[…]
@@
where $A_{j-1}=D_{j-1}W^{r}-I$ and $G_{j-1}$ is a rank-1 matrix.
Expanding the matrix product to the first order in the rank-1 corrections $G_{j-1}$, we obtain
@@
The first term describes the main dynamics without the rank-1 updates, while the second term is a sum of rank-1 corrections, each inserted at a different time step $m$ and propagated by the main dynamics. We now factor the gate scalars $g_{j-1}$ explicitly from the first term in (47) like we did in Section 5.2.1.
@@
$\mu^{*}_{t,k}\;=\;\mu\prod_{j=k+1}^{t}g_{j-1}\;+\;\text{(perturbative correction terms)}.$ (50)
Here the perturbative correction terms arise from the residual $(1-g_{t})x_{t}$ pathway and from gate-gradient contributions. These introduce rank-1 modifications to the Jacobian products, modulating gradient propagation in a way that cannot be collapsed into a single scalar factor. Thus, while the multiplicative gate product sets the dominant time-varying decay, the perturbative corrections adjust the effective learning rate in a lag- and state-dependent manner.
@@ 5.2.3 RNNs with multiple gates
The first term describes the main dynamics through diagonal gates and normalized matrices $\tilde{A}_{j-1}$. The second term accounts for corrections due to $G_{j-1}$.
@@
$\mu^{*(i)}_{t,k}\;=\;\mu\prod_{j=k+1}^{t}g^{(i)}_{j-1}\;+\;\text{(perturbative correction terms)},$ (58)
where $\mu^{*(i)}_{t,k}$ is the effective learning rate associated with neuron $i$.
Here the perturbative correction terms involve interactions between gates and state updates, producing full-rank contributions that cannot be simplified into a scalar or diagonal factor. Their effect is modulated by the diagonal gate products but introduces additional couplings across neurons. In the multi-gate case:
  • Each neuron has its own time-varying effective learning rate determined by its specific gate trajectory.
  • Gradient propagation is anisotropic: some directions are damped strongly, while others are preserved, depending on the distribution of gate trajectories across neurons.
  • The correction terms act as neuron-coupling perturbations, making the effective scaling both magnitude- and direction-dependent.
6 Analogies between gating mechanisms and adaptive gradient descent methods
@@
With a scalar gate $g_{j-1}$, the gradient is modulated by a global, input-driven attenuation factor that varies over time. This is conceptually similar to a learning rate schedule (e.g., exponential decay), with the key difference that the modulation emerges from the network’s own state dynamics rather than from an externally prescribed schedule.
@@
In the multi-gate case, the modulation is neuron-specific, so each neuron $i$ has its own effective learning rate, determined by the trajectory of its gate values. This mirrors adaptive optimizers such as Adam or RMSProp, which assign each parameter a distinct, dynamically adjusted step size.